### v2

@@ 4.3 Potential Bias of Manually Defining an EHR Text Serialization
Defining an EHR serialization involved subjective decisions, which may have introduced bias. For instance, awareness of the prediction tasks could influence the prioritization of certain data elements, potentially favoring task-relevant information. To minimize this risk, we aimed to create an objectively defined serialization that encapsulates key aspects of the EHR records. Also, due to computational constraints, we evaluated only three variants of our final serialization for 4,096 tokens: (1) appending a list of all unique conditions at the beginning, (2) omitting the three values for the listed comments (e.g., for “Cigarette consumption” and “pH measurement, venous” in Fig. 2), and (3) combining both approaches. For GTE-Qwen2-7B, the chosen serialization performed best, but all variants outperformed CLMBR-T-Base. However, for LLM2Vec-Llama-3.1-8B, only the serialization omitting all values (variant 2) performed slightly better. We selected the current serialization for its simplicity and completeness of medical information. This approach avoided introducing additional entries of all unique conditions and preserved the most recent values for visit-level concepts.
4.4 LLM-Embedding Models and Baselines
In this study, we evaluated two LLM-embedding models, GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B, based on state-of-the-art decoder-only LLMs. These models were selected for their ability to handle the 4.096-token EHR serializations used in our experiments. For comparison, we also tested a smaller variant of both models. As additional baselines we included commonly used encoder-only embedding models with smaller input sizes (512 tokens). To use them with 4,096 token inputs, the EHR serializations were split into up to eight 512-token chunks, and the resulting embeddings were averaged to generate a single representation. For all LLM-embedding models and smaller language models used in this study, we used the mean pooling of the last layer as the final embedding [26, 44]. The LLM2Vec models used a slight variation that only incorporates the tokens that do not belong to the instruction. Below is an overview of all models:
@@ 4.4 LLM-Embedding Models and Baselines
This LLM-embedding model is based on the Qwen2-7B-Instruct LLM [31] using a decoder-only Transformer architecture with 28 layers, 28 attention heads, and a hidden size of 3.584444Hugging Face identifier: Alibaba-NLP/gte-Qwen2-7B-instruct. It was trained with autoregressive next token prediction and converted into an embedding model using the General Text Embedding (GTE) method [29]. This conversion replaces causal attention with bidirectional attention, enabling the model to attend to both left and right contexts for token embedding, akin to BERT. Contrastive learning was applied using a mixture of private datasets to enhance embedding performance. The model also incorporates instructions tailored for embedding tasks, supporting a context size of up to 32.000 tokens.
@@
A smaller variant of GTE-Qwen2-7B, this model is based on Qwen2-1.5B-Instruct, with 28 layers, 12 attention heads, and a hidden size of 1,536. It was also trained using the GTE method [29] and supports a context size of up to 32.000 tokens.
@@
This model is built upon the Llama-3.1-8B-Instruct LLM [32] with a decoder-only Transformer architecture with 32 layers, 32 attention heads, and a hidden size of 4.096555Hugging Face identifier: McGill-NLP/LLM2Vec-Meta-Llama-31-8B-Instruct-mntp-supervised. Initially trained for next-token prediction, it was converted to an embedding model using the LLM2Vec method [26]. This method adds bidirectional attention and fine-tunes the model with supervised contrastive learning on embedding tasks. The finetuning used curated data from the public E5 dataset [45, 46], containing approximately 1.5 million entries. The model supports task-specific instructions and accommodates a context size of up to 128.000 tokens.
@@
DeBERTa v3 is an encoder-only Transformer model designed for token embeddings [48]. It improves upon its predecessor by replacing the masked language modeling objective with replaced token detection and utilizing Gradient-Disentangled Embedding Sharing. We evaluated the base variant (12 layers, 12 attention heads, 768 hidden size) and the large variant (24 layers, 12 attention heads, 1,024 hidden size), with parameter counts of 183M and 434M, respectively666Hugging Face identifiers: microsoft/deberta-v3-base,large.
@@
BERT is a well-established text embedding model using an encoder-only transformer trained with the masked language modelling objective [15]. We included both the base (12 layers, 12 attention heads, 768 hidden size, 110M parameters) and large (24 layers, 16 attention heads, 1,024 hidden size, 340M parameters) variants as benchmarks777Hugging Face identifiers: google-bert/bert-base,large-uncased. While not state-of-the-art, BERT models remain widely used in embedding tasks.
@@
CLIMBR-T-Base
CLMBR-T-Base is a specialized EHR foundation model trained on 2.57 million deidentified EHRs from Stanford Medicine with autoregressive next code prediction [13, 30]. It uses gated recurrent units and has 12 layers and a hidden dimension of 768999Hugging Face identifier: StanfordShahLab/clmbr-t-base. The model has 141M parameters and allows for a context window of 496 codes. CLIMBR-T-Base has demonstrated consistent improvements over count-based baselines for a variety of clinical prediction tasks [30]. It serves as a main baseline for our experiments to test specialized EHR models against general purpose text embeddings models for representing EHR records.
LLM-Embedding Model and CLIMBR-T-Base
To test whether the LLM-embedding models and the EHR foundation model learn orthogonal information, we combined both models for the prediction. To this end we simply appended both embeddings. The resulting embeddings have dimensions 4.352 (GTE-Qwen2-7B) and 4.864 (LLM2Vec-Llama-3.1-8B).
Counts-based Model
Counts models have proven to be strong baselines for EHR prediction tasks [6, 12, 13]. The basic idea is to encode all EHR events of a patient in a single vector where each entry represents the number of occurrences of a medical concept. We used the counts baseline introduced in [30] that further extends this approach with ontology expansion, enriching the vectors with parent and child concepts.
Based on the embeddings or the counts vectors given by the methods described above a classification head was trained and validated for each prediction task. For the embedding models we used a logistic regression head. For the counts-based model, we used a gradient boosting machine [39], which proved superior [30]. We adopted the parameter tuning of the classification heads from the EHRSHOT benchmark to ensure comparability of results.
4.5 Instructions for LLM-Embedding Models
The GTE and LLM2Vec models used instruction-tuned embeddings, requiring task-specific prompts. Hence, we added simple instructions for each prediction task based on their respective instruction templates. For instance, for prediction of anemia we added “Given a patient’s electronic healthcare record (EHR) in Markdown format, retrieve relevant passages that answer the query: has the patient anemia”. The existing EHRSHOT benchmark encoded the EHRs of the same patient and the identical prediction times only once for efficiency reasons. However, to support task-specific instructions, we had to change this default behavior leading to 1.161.412 instead of 406.379 EHRs to encode resulting in longer processing times. The difference between 1.161.412 labels used in our experiments and the total number of labels of 1.178.665 (Table 2) is since some labels even have the same tasks and prediction time and are merged. We list all instructions in Table 8 and perform ablations testing the effect of the instructions.
@@ 4.6 Computational Setup and Running Times
All experiments were conducted on the Charité High-Performance Cluster using Nvidia A100 GPUs with 80 GB memory, configured with one, four, or eight GPUs (DGX systems). Running the GTE-Qwen2-7B model with our serialization truncated at 4,096-tokens on an 8-GPU DGX system required approximately 20 hours. For the LLM2Vec-Llama-3.1-8B model, runtime errors occurred during multi-GPU experiments with the full dataset. These issues were resolved by splitting the data into smaller batches, which introduced additional overhead. Additionally, we optimized the LLM2Vec code by removing an initial text pruning step that took approximately eight hours when using the full data. Using this modified setup, the LLM2Vec-Llama-3.1-8B model took approximately 30 hours to complete on eight A100 GPUs.
4.7 Performance Results on 15 Prediction Tasks and Few-Shot Setting
@@
Following the EHRSHOT benchmark, we evaluated all models across 15 prediction tasks under various few-shot settings. The benchmark includes a modular pipeline designed to execute key tasks, with the flexibility to optionally utilize a Slurm cluster for distributed execution. Running all steps within this pipeline ensures full reproducibility of results. Step four of the pipeline, which generates EHR representations with CLIMBR-T-Base and the counts-based model, was extended to incorporate our method for creating language model-based EHR representations. This adaptation allowed us to reuse significant portions of the existing code, including the task evaluation framework. Additionally, we implemented new functionality for EHR serialization and slightly modified other steps of the benchmark to accommodate our experimental setup. For instance, the label creation process was adjusted (step three) to enable task-specific instructions for the LLM-embedding models. All modifications have been documented and can be tracked in our public GitHub repository.
4.8 Effect of Different Contexts Sizes
@@
We investigated the impact of varying context sizes in the LLM-embedding models. We wanted to determine whether encoding information from older visits enhances prediction performance and whether longer inputs might dilute critical details, such as laboratory values, in the final embeddings. Specifically, we evaluated GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B models with input token limits of 512, 1.024, 2.048, and 8.192 tokens. Input tokens exceeding these thresholds were discarded. Due to the design of our EHR serialization process, additional input tokens primarily consisted of medical concepts from past visits. By testing these varying context sizes, we aimed to assess the balance between capturing historical medical data and preserving the clarity of high-priority information within the embeddings.
@@ 4.9 Effect of Chunked Contexts
To further explore whether LLM-embedding models effectively interpret the entirety of their input, we compared the performance of models processing complete inputs versus segmented (chunked) inputs. For this, the 4.096-token inputs were divided into smaller chunks of sizes 512, 1.024, and 2.048 tokens101010In our implementation we used these chunks sizes decreased by eight tokens to cater for special tokens that might be added by the implementations of the language models.. Separate embeddings were generated for each chunk, and a final embedding was obtained by averaging the embeddings of all chunks. This approach aligns with the mean pooling applied to the last layer that we used to compute embeddings. In this setup, task-specific instructions, placed at the beginning of the inputs, were included only in the first chunk. This design allowed us to evaluate the models’ ability to contextualize task instructions and whether chunking affects the overall performance.
4.10 Ablations of EHR Serialization
To better understand the contribution of various components in the EHR serialization process to the performance of the LLM-embedding models, we conducted a series of ablation experiments. First, we assessed the impact of task-specific instructions by removing them entirely to determine their influence on the final embeddings. Subsequently, we performed additional ablations by systematically excluding specific components of the serialization. These included demographic data, aggregated LOINC codes, and both visit summaries and detailed entries. Furthermore, we examined the effect of removing specific fields from the detailed visit entries, such as conditions, medications, and procedures. Throughout these experiments, the rest of the pipeline was kept consistent to isolate the effects of the removed components. This approach allowed us to identify which parts of the serialization process were most critical for generating effective EHR representations, providing insights into how these models leverage structured medical data for prediction tasks.

### v3

@@ 4.3 Potential Bias of Manually Defining an EHR Text Serialization
Defining an EHR serialization involved subjective decisions, which may have introduced bias. For instance, awareness of the prediction tasks could influence the prioritization of certain data elements, potentially favoring task-relevant information. To minimize this risk, we aimed to create an objectively defined serialization that encapsulates key aspects of the EHR records. Also, due to computational constraints, we evaluated only three alternatives to our final serialization for 4,096 tokens: (1) appending a list of all unique conditions at the beginning, (2) omitting the three values for the listed comments (e.g., for “Cigarette consumption” and “pH measurement, venous” in Fig. 2), and (3) combining both approaches. For GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B, the EHR serialization chosen for our experiments performed on par or better than the three alternatives. We also selected the current serialization for its simplicity and completeness of medical information. This approach avoided introducing additional entries of all unique conditions and preserved the most recent values for visit-level concepts.
4.4 LLM Embedding Models and Baselines
In this study, we evaluated two LLM embedding models, GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B, based on state-of-the-art decoder-only LLMs. We selected these models for their ability to handle the 4,096-token EHR serializations used in our experiments. For comparison, we also tested a smaller variant of both models. As additional baselines, we included commonly used encoder-only embedding models with smaller input sizes (512 tokens). To use them with 4,096 token inputs, the EHR serializations were split into up to eight 512-token chunks, and the resulting embeddings were averaged to generate a single representation. For all LLM embedding and smaller language models, we used mean pooling of the last layer for the final embedding [27, 46]. The LLM2Vec models used a slight variation that only incorporates the tokens that do not belong to the instruction. Below is an overview of all models:
@@ 4.4 LLM Embedding Models and Baselines
This LLM embedding model is based on the Qwen2-7B-Instruct LLM [32], which uses a decoder-only Transformer architecture with 28 layers, 28 attention heads, and a hidden size of 3,584444Hugging Face identifier: Alibaba-NLP/gte-Qwen2-7B-instruct. It was trained with autoregressive next token prediction and converted into an embedding model using the General Text Embedding (GTE) method [30]. This conversion replaces causal attention with bidirectional attention, enabling the model to attend to both left and right contexts for token embedding, akin to BERT. Contrastive learning was applied using a mixture of private datasets to enhance embedding performance. The model also incorporates instructions tailored for embedding tasks and supports a context size of up to 32,000 tokens.
@@
A smaller variant of GTE-Qwen2-7B, this model is based on Qwen2-1.5B-Instruct, with 28 layers, 12 attention heads, and a hidden size of 1,536. It was also trained using the GTE method [30] and supports a context size of up to 32,000 tokens.
@@
This model is built upon the Llama-3.1-8B-Instruct LLM [33], which has a decoder-only Transformer architecture with 32 layers, 32 attention heads, and a hidden size of 4,096555Hugging Face identifier: McGill-NLP/LLM2Vec-Meta-Llama-31-8B-Instruct-mntp-supervised. Initially trained for next-token prediction, it was converted to an embedding model using the LLM2Vec method [27]. This method adds bidirectional attention and fine-tunes the model with supervised contrastive learning on embedding tasks. The fine-tuning used curated data from the public E5 dataset [47, 48], containing approximately 1.5 million entries. The model supports task-specific instructions and supports a context size of up to 128,000 tokens.
@@
DeBERTa v3 is an encoder-only Transformer model designed for token embeddings [50]. It improves upon its predecessor by replacing the masked language modeling objective with replaced token detection and using Gradient-Disentangled Embedding Sharing. We evaluated the base variant (12 layers, 12 attention heads, 768 hidden size) and the large variant (24 layers, 12 attention heads, 1,024 hidden size), with parameter counts of 183M and 434M, respectively666Hugging Face identifiers: microsoft/deberta-v3-base,large.
@@
BERT is a well-established text embedding model using an encoder-only Transformer trained with the masked language modeling objective [15]. We included both the base (12 layers, 12 attention heads, 768 hidden size, 110M parameters) and large (24 layers, 16 attention heads, 1,024 hidden size, 340M parameters) variants as benchmarks777Hugging Face identifiers: google-bert/bert-base,large-uncased. While not state-of-the-art, BERT models remain widely used in embedding tasks.
@@
CLMBR-T-Base
CLMBR-T-Base is a specialized EHR foundation model trained on 2.57 million de-identified EHRs from Stanford Medicine with autoregressive next code prediction [13, 31]. It uses gated recurrent units and has 12 layers and a hidden dimension of 768999Hugging Face identifier: StanfordShahLab/clmbr-t-base. The model has 141M parameters and allows for a context window of 496 codes. CLMBR-T-Base has demonstrated consistent improvements over count-based baselines for a variety of clinical prediction tasks [31]. It serves as a main baseline for our experiments to test specialized EHR models against general-purpose text embedding models for representing EHR records.
LLM Embedding Model and CLMBR-T-Base
To test whether the LLM embedding models and the EHR foundation model learn orthogonal information, we combined both models for the prediction. To this end, we simply appended both embeddings. The resulting embeddings have dimensions 4,352 (GTE-Qwen2-7B) and 4,864 (LLM2Vec-Llama-3.1-8B).
Count-based Model
Count models have proven to be strong baselines for EHR prediction tasks [6, 12, 13]. The basic idea is to encode all EHR events of a patient in a single vector where each entry represents the number of occurrences of a medical concept. We used the counts baseline introduced in [31] that further extends this approach with ontology expansion, enriching the vectors with parent and child concepts.
Based on the embeddings or the counts vectors generated by the methods described above, a classification head was trained and validated for each prediction task. For the embedding models we used a logistic regression head. For the count-based model, we used a GBM [41], which proved superior [31]. We adopted the parameter tuning of the classification heads from the EHRSHOT benchmark to ensure comparability of results.
4.5 Instructions for LLM Embedding Models
The GTE and LLM2Vec models used instruction-tuned embeddings, requiring task-specific prompts. Hence, we added simple instructions for each prediction task based on their respective instruction templates. For instance, for prediction of anemia, we added “Given a patient’s electronic healthcare record (EHR) in Markdown format, retrieve relevant passages that answer the query: has the patient anemia”. The existing EHRSHOT benchmark encoded the EHRs of the same patient and the identical prediction times only once for efficiency reasons. However, to support task-specific instructions, we had to change this default behavior, leading to 1,161,412 instead of 406,379 EHRs to encode, resulting in longer processing times. The difference between 1,161,412 labels used in our experiments and the total number of labels of 1,178,665 (Table 2) is because some labels share the same task and prediction time and are therefore merged. For assignment of new diagnoses tasks in UKB, the same prompt was used for all diseases to reduce computational resources. We list all instructions in Table 8 and perform ablations to test the effect of the instructions.
@@ 4.6 Computational Setup and Running Times
All experiments were conducted on the Charité High-Performance Cluster using Nvidia A100 GPUs with 80 GB memory, configured with one, four, or eight GPUs (DGX systems). Running the GTE-Qwen2-7B model with our serialization truncated at 4,096 tokens on an 8-GPU DGX system required approximately 20 hours. For the LLM2Vec-Llama-3.1-8B model, runtime errors occurred during multi-GPU experiments with the full dataset. These issues were resolved by splitting the data into smaller batches, which introduced additional overhead. Additionally, we optimized the LLM2Vec code by removing an initial text pruning step that took approximately eight hours when using the full data. Using this modified setup, the LLM2Vec-Llama-3.1-8B model took approximately 30 hours to complete on eight A100 GPUs. The calculation of the embeddings for the 387,464 UKB patients took approximately 35 hours on a single GPU per task for the LLM-based methods. For CLMBR-T-Base, the calculation was faster, only requiring around eight hours.
4.7 Performance Results on EHRSHOT Prediction Tasks and Few-Shot Setting
@@
Following the EHRSHOT benchmark, we evaluated all models across 15 prediction tasks under various few-shot settings. The benchmark includes a modular pipeline designed to execute key tasks, with the flexibility to optionally use a Slurm cluster for distributed execution. Running all steps within this pipeline ensures full reproducibility of results. Step four of the pipeline, which generates EHR representations with CLMBR-T-Base and the count-based model, was extended to incorporate our method for creating language model-based EHR representations. This adaptation allowed us to reuse significant portions of the existing code, including the task evaluation framework. Additionally, we implemented new functionality for EHR serialization and slightly modified other steps of the benchmark to accommodate our experimental setup. For instance, the label creation process was adjusted (step three) to enable task-specific instructions for the LLM embedding models. All modifications have been documented and can be tracked in our public GitHub repository.
4.8 External Validation on UK Biobank
External validation was performed using data from the UKB, a large-scale prospective cohort study comprising 502,489 UK participants recruited between 2006 and 2010, with a median follow-up of 13.8 years. We used linked EHR data from primary care (General Practitioner, GP) and secondary care (Hospital Episode Statistics, HES), providing information on diagnoses, procedures, and prescriptions.
Initial data preprocessing, including cleaning, feature extraction, missing value imputation, and endpoint selection followed the methodology described in [34]. All health records were mapped to the OMOP CDM using mapping tables provided by the UKB, SNOMED International, and the OHDSI community for mapping concepts from the provider and country-specific non-standard vocabularies to OMOP standard vocabularies. Participants lacking any recorded GP or HES events either before or after their recruitment date were excluded, resulting in a validation cohort of 387,464 individuals. Diagnostic codes were mapped to Phecodes X [53, 54] (derived directly from source ICD-10 or by mapping from SNOMED to ICD-10 codes and subsequently to Phecodes X) primarily for standardized endpoint definition and cohort selection. To avoid redundancy with source codes used as features, the Phecodes themselves were excluded during the creation of patient sequences for model input. Due to significant challenges in mapping and harmonizing UKB laboratory values [55], and to ensure comparability across models and tasks, laboratory data were excluded entirely, differing from the use of binary labels in [34]. The final feature set comprised conditions (SNOMED, CVX), medications (RxNorm), and procedures (SNOMED).
Given that UKB represents a general population cohort, we adjusted the proposed prediction tasks to define relevant longitudinal health trajectory tasks. These tasks included: (1) prediction of all-cause hospitalization within the next year (operational outcomes), (2) prediction of incident diagnoses for a set of selected conditions (assignment of new diagnoses), and (3) prediction of all-cause mortality (mortality prediction). The selection of incident diseases for the assignment of new diagnoses task group largely followed [34], focusing on common conditions, diseases lacking established risk stratification tools, and specific cardiovascular conditions. From the initial 24 endpoints proposed in [34], we treated all-cause death as a separate task. For the assignment of new diagnoses and mortality tasks, patients with a diagnosis of the respective endpoint recorded prior to their UKB recruitment date were excluded to prevent data leakage; this exclusion was not applied to the hospitalization task due to high incidence. Following task-specific exclusions, the final sample size available for analysis ranged from 300,344 to 359,250 patients, depending on the prediction task (see Table 6 for details). For all three task groups, the prediction window was set to one year after the prediction year.
The pre-trained CLMBR-T-Base model operates with a fixed vocabulary (65,536 total, 26,249 unique codes). To use this model, we mapped the 50,702 unique medical codes (SNOMED CT, RxNorm, CVX) present in our processed UKB cohort to the CLMBR-T-Base vocabulary. This mapping followed the steps similar to those implemented in the FEMR package101010https://github.com/som-shahlab/femr/tree/0ebf454303091fa83c9e4023e7db9ba9b68af09a, involving direct code matching where possible, supplemented by indirect mapping via the OHDSI ATHENA vocabulary using ’Maps to’ relationships and inclusion of ancestor concepts. This process successfully mapped 7,969 unique UKB codes to the CLMBR-T-Base vocabulary (in format ontology/code), which means that the remaining codes were implicitly excluded from the input sequences for this model. After creating patient timelines, adding information about birth date, ethnicity, and visits, the data was converted into the MEDS standard [56], from which the embeddings were calculated. Additionally, UKB ethnicities were converted to the ethnicity groups used by CLMBR-T-Base. For transforming the patient information into embeddings, we mainly followed the code provided by FEMR (mainly the convert_patient function111111https://huggingface.co/StanfordShahLab/clmbr-t-base) with minor modifications to enable batch processing.
Similar to the approach of Wornow et al. [31], we performed an ontology expansion to improve the counts baseline by adding direct parent and grandparent concepts. This increased the unique code count from 51,677 to 69,850, aiming to capture broader hierarchical relationships between medical concepts. However, the resulting high dimensionality posed memory challenges for a standard counts matrix. To mitigate this, we implemented feature selection by removing codes present in less than 50 individuals (0.01% of the cohort), reducing the feature space to 24,509 unique codes. To ensure a more comparable baseline to the other models, which implicitly incorporate patient context, we explicitly added normalized age at prediction time and coded sex as additional features to the feature matrix used for the count-based baseline.
Similar to before, all approaches were evaluated using a five-fold cross-validation. Within each fold, hyperparameters were optimized on a dedicated validation set of similar sample size to that of the training partition in that fold. For each fold, patients were randomly assigned to train, validation, and test sets. The test set for each task within a fold was constructed by selecting all available positive cases (capped to 10,000), and an equal number of negative cases after allocating patients to the train/validation sets. To specifically assess performance in low-data scenarios, mimicking potential clinical applications with limited examples, we again evaluated models using incrementally larger subsets of the training data, with the final evaluation point capped at a sample size of 256 patients from the train set, even when larger test sets were available.
4.9 Ablations of EHR Serialization
To better understand the contribution of various components in the EHR serialization process to the performance of the LLM embedding models, we conducted a series of ablation experiments. First, we assessed the impact of task-specific instructions by replacing them with a generic instruction (Table 8) or removing them entirely to determine their influence on the final embeddings. Subsequently, we performed additional ablations by systematically excluding specific components of the serialization. These included demographic data, aggregated LOINC codes, and both visit summaries and detailed entries. Furthermore, we examined the effect of removing specific fields from the detailed visit entries, such as conditions, medications, and procedures. Throughout these experiments, the rest of the pipeline was kept consistent to isolate the effects of the removed components. This approach allowed us to identify which parts of the serialization process were most critical for generating effective EHR representations, providing insights into how these models use structured medical data for prediction tasks.
4.10 Effect of Different Time Windows
To examine the influence of recency on predictive performance, we varied the time window preceding the prediction date used during EHR text serialization. We evaluated GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B across six intervals: one day, one week (7 days), one month (30 days), one year (365 days), three years (1,095 days), and full history. For each window, only events occurring within the specified interval before the prediction time were included. Hence, only data from the respective time window contributed to the aggregated information and the visit data in the EHR serialization. All other aspects of the serialization, including structure, formatting, and instruction prompts remained unchanged to isolate the effect of the temporal window.
4.11 Effect of Different Context Sizes
@@
We investigated the impact of varying context sizes in the LLM embedding models. We wanted to determine whether encoding information from older visits enhances prediction performance and whether longer inputs might dilute critical details, such as laboratory values, in the final embeddings. Specifically, we evaluated GTE-Qwen2-7B and LLM2Vec-Llama-3.1-8B models with input token limits of 512, 1,024, 2,048, and 8,192 tokens. Input tokens exceeding these thresholds were discarded. Due to the design of our EHR serialization process, additional input tokens primarily consisted of medical concepts from past visits. By testing these varying context sizes, we aimed to assess the balance between capturing historical medical data and preserving the clarity of high-priority information within the embeddings.
@@ 4.12 Effect of Chunked Contexts
To further explore whether LLM embedding models effectively interpret their entire input, we compared the performance of models processing complete inputs versus segmented (chunked) inputs. For this, the 4,096-token inputs were divided into smaller chunks of sizes 512, 1,024, and 2,048 tokens121212In our implementation, we used these chunk sizes decreased by eight tokens to cater for special tokens that might be added by the implementations of the language models.. Separate embeddings were generated for each chunk, and a final embedding was obtained by averaging the embeddings of all chunks. This approach aligns with the mean pooling applied to the last layer that we used to compute embeddings. In this setup, task-specific instructions, placed at the beginning of the inputs, were included only in the first chunk. This design allowed us to evaluate the models’ ability to contextualize task instructions and whether chunking affects the overall performance.
Acknowledgements
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – Project-ID 437531118 – SFB 1470 and German Federal Ministry of Education and Research (BMBF) Project-ID 01ZZ2317G.
Declaration of interest
The authors declare no competing interests.
Data availability
The EHRSHOT data is available through gated access via https://doi.org/10.57761/0gv9-nd83. UK Biobank data, including all linked routine health records, are publicly available to bona fide researchers upon application at http://www.ukbiobank.ac.uk/using-the-resource. In this study, only primary care data not subject to the Government’s Control of Patient Information (COPI) notice was used (UK Biobank Category 3000).
Code availability
All extensions and experiments conducted for this paper have been made publicly available via our GitHub repository: https://github.com/stefanhgm/ehrshot-benchmark.